### v1

# Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory

Abstract
The evaluation of large language models (LLMs) via benchmarks is widespread, yet inconsistencies between different leaderboards and poor separability among top models raise concerns about their ability to accurately reflect authentic model capabilities. This paper provides a critical analysis of benchmark effectiveness, examining main-stream prominent LLM benchmarks using results from diverse models. We first propose a new framework for accurate and reliable estimations of item characteristics and model abilities. Specifically, we propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture. Based on PSN-IRT, we conduct extensive analysis which reveals significant and varied shortcomings in the measurement quality of current benchmarks. Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference [^1] Code and data are openly available at https://github.com/Joe-Hall-Lee/PSN-IRT.
Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory
[…]
footnotetext: * The first two authors contributed equally.
1 Introduction
Figure 1: Illustration of weak separability and ranking inconsistencies in LLM benchmarks [^3] https://opencompass.org.cn.
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important. Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions aimed at assessing model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting deeper consideration of their effectiveness McIntosh et al. (2024).
As illustrated in Figure 1, on one hand, different benchmarks often produce inconsistent leaderboards Perlitz et al. (2024), with the same model achieving substantially varying rankings across evaluations. This makes it challenging for researchers and developers to determine a model’s real performance. On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^4] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets and its assumption of normally distributed abilities does not always hold.
To advance our objective, we propose the Pseudo-Siamese Network for IRT (PSN-IRT), a framework for comprehensive and interpretable analysis of benchmark test sets. PSN-IRT processes model identifiers and item identifiers through independent neural network pathways. These pathways respectively estimate latent model ability and a rich set of item parameters. The estimated properties are then integrated using an IRT-based formulation to predict outcome probabilities. This architectural design ensures both powerful modeling capabilities and strong theoretical interpretability.
To validate the effectiveness of PSN-IRT, we conduct extensive experiments on evaluation results from 12 LLMs across 11 benchmarks. Our results show that PSN-IRT outperforms previous approaches in both parameter estimation accuracy and reliability. Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with key findings as follows:
  1. 1.
    LLM benchmarks fail to achieve simultaneous excellence across multiple measurements.
  2. 2.
    LLM benchmarks suffer from widespread saturation and insufficient difficulty ceilings, limiting their ability to challenge and accurately evaluate advanced models.
  3. 3.
    LLM benchmarks exhibit data contamination in numerous items, compromising their reliability.
Furthermore, we find that model rankings derived from high-quality datasets selected by PSN-IRT are more consistent with human preference and offer stronger discriminability among top models.
Our contributions are summarized as follows:
  1. 1.
    We propose PSN-IRT, a benchmark analysis framework demonstrating superior estimation accuracy and reliability.
  2. 2.
    We conduct extensive experiments, using PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in may aspects.
  3. 3.
    We show that selecting items with PSN-IRT leads to model rankings that better align with human preference.
2 Background
2.1 Related Work
Researchers have explored various methods to enhance the reliability and interpretability of LLM benchmark evaluations. During test set construction, many benchmarks incorporate human annotation and leverage strong LLMs to ensure item quality, detect inconsistencies, and manage item difficulty He et al. (2024); Wei et al. (2024). Beyond these construction-phase strategies, recent studies have also introduced quantitative metrics to assess benchmark behavior post-deployment. For example, Benchmark Agreement Testing (BAT) evaluates the reliability of new benchmarks by comparing them with established ones using correlation-based measures Perlitz et al. (2024); White et al. (2025). Other work proposes novel metrics like Separability with Confidence for more nuanced benchmark analysis Li et al. (2024).
Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004) offers a psychometric framework for finer-grained, item-level analysis. Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability Byrd and Srivastava (2022); Vania et al. (2021) and, more recently, to inspire methods for improving evaluation efficiency, for instance, through optimized item selection or adaptive testing paradigms Maia Polo et al. (2024); Zhuang et al. (2024). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored.
Figure 2: The illustration of our proposed PSN-IRT. Separate neural networks estimate model ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response. After that, the networks can be leveraged for estimating properties for models or items, respectively.
2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) Lord et al. (1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) DeVellis (2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
Central to IRT is the Item Characteristic Curve (ICC), a mathematical function that describes the probability of a correct response, $P(X=1\mid\theta)$, as a function of ability $\theta$. Typically logistic, the ICC visualizes how item properties influence performance, serving as the foundation for IRT.
The simplest IRT model, the One-Parameter Logistic (1PL) model, defines the ICC as:
$P(X=1\mid\theta)=\frac{1}{1+e^{-(\theta-b)}}$ (1)
where $b$ denotes item difficulty. When $\theta=b$, the probability of the examinee generating a correct response is 0.5. The 1PL assumes all items have equal discriminability, a simplification that may not hold in complex testing scenarios.
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data Tsutsumi et al. (2021).
3 Pseudo-Siamese Network for IRT
In this section, we propose the Pseudo-Siamese Network for IRT (PSN-IRT), designed to diagnose benchmark item quality by analyzing LLM responses. Specifically, one property can be inferred for each model: model-ability, and four properties can be inferred for each benchmark item: discriminability, difficulty, guessing-rate and feasibility.
Model Architecture.
The PSN-IRT architecture comprises two independent networks: a model network and an item network. The model network processes one-hot encoded LLM identifiers to estimate model-ability $\theta$, while the item network handles one-hot encoded item identifiers to produce the four IRT parameters: discriminability $a$, difficulty $b$, guessing-rate $c$, and feasibility $d$.
Each network consists of three fully connected layers with ReLU activation functions, enabling efficient and stable parameter estimation. After that, these estimated properties are then processed by a Logistic Calculation Layer. This layer operates using the Four-Parameter Logistic (4PL) model, an advanced IRT formulation that extends simpler models like the 1PL to capture nuanced behaviors:
$P(X=1\mid\theta)=c+(d-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (2)
where each parameter controls a different aspect of benchmark item behavior:
  • The difficulty $b$ determines the ability level where the probability changes most significantly.
  • The discriminability $a$ controls the steepness.
  • The guessing-rate $c$ captures how likely models are to succeed without full understanding.
  • The feasibility $d$ represents the maximum probability that even highly proficient models will correctly answer the item.
Training and Applications.
The PSN-IRT is trained end-to-end using the observed binary response data in the form of (Model, Item, Response, Outcome), where the binary outcome indicates the correctness of each LLM on each benchmark item. During training , PSN-IRT processes encoded pairs of LLM and item identifiers through their respective network pathways. For each pair, the PSN-IRT framework first internally estimates the specific model and item properties. These estimated properties are then passed to a Logistic Calculation Layer, with the training objective as follows:
$J(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(P_{i})+(1-y_{i})\log(1-P_{i})\right]$ (3)
where $N$ is the batch size. The training objective is to minimize the discrepancy between the model’s prediction and the observed binary outcomes. Consequently, all learnable weights within both the model and item networks are updated simultaneously, concurrently optimizing the estimated properties for both models and benchmark items.
Upon completion of training, the two networks of PSN-IRT can be used for evaluating model performance and diagnosing benchmark characteristics, respectively. Specifically, the Model Network, by processing a model’s encoding, outputs an estimate of the model ability $\theta$. Meanwhile, the Item Network, given an benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
4 Experiments
4.1 Set-up
Dataset.
To evaluate existing LLM benchmarks, we select 11 datasets that vary in domain coverage. They serve as a basis for comparing different IRT-based methods and analyzing test set properties. An overview of the datasets is provided in Table 1.
To construct training data, we conduct evaluations using OpenCompass Contributors (2023). The evaluation results are collected in the form of binary outcomes, indicating whether each model answered each item correctly. Then, we construct a unified outcome matrix by aggregating item-level results across all models and benchmarks. We split the interactions into training (60%), validation (20%), and test sets (20%). The validation set is utilized to monitor performance during model training and to implement early stopping, thereby preventing overfitting.
Benchmark Data Size Domain Metric Format
ARC-C Clark et al. (2018) 295 General EM Multiple Choice
BBH Suzgun et al. (2023) 6,511 General EM Mixed
Chinese SimpleQA He et al. (2024) 3,000 Knowledge LLM-as-a-Judge QA
GPQA Diamond Rein et al. (2024) 198 Science EM Multiple Choice
GSM8K Cobbe et al. (2021) 1,319 Math EM QA
HellaSwag Zellers et al. (2019) 10,042 General EM Multiple Choice
HumanEval Chen et al. (2021) 164 Code Pass@1 QA
MATH Hendrycks et al. (2021b) 5,000 Math EM QA
MBPP Austin et al. (2021) 500 Code Pass@1 QA
MMLU Hendrycks et al. (2021a) 14,042 General EM Multiple Choice
TheoremQA Chen et al. (2023) 800 Science EM QA
Table 1: Benchmarks used in this paper.
Model Parameter Method ACC F1 AUC Kendall Average
IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 0.7070 0.7811 0.7278 0.9697 0.7964
VI 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT 1PL Deep Learning 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT 4PL Deep Learning 0.7991 0.8520 0.8477 1.0000 0.8747
Table 2: Comparison of prediction accuracy and rank reliability across different methods.
Model.
We mainly focus on currently high-performing models, including 360GPT2-Pro [^5] https://ai.360.cn/, DeepSeek-V3 DeepSeek-AI et al. (2025), Doubao-Pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5 Team et al. (2024a), Hunyuan-Turbo [^7] https://hunyuan.tencent.com/, Moonshot-v1888https://platform.moonshot.cn/, Qwen-Plus Yang et al. (2024), and Yi-Lightning Wake et al. (2025). However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality Martínez-Plumed et al. (2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it Team et al. (2024b), Mistral-7B-Instruct-v0.1 Jiang et al. (2023), Qwen2.5-3B-Instruct Yang et al. (2024), and Vicuna-7B-v1.3 Chiang et al. (2023).
Baseline.
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) Hastings (1970), Variational Inference (VI) Jordan et al. (1999), and VIBO Wu et al. (2020). We also compare with Deep-IRT Tsutsumi et al. (2021), which is a neural extension of 1PL IRT that learns item-trait interactions via deep networks.
Metric.
Two quantitative metrics are included:
  • Prediction accuracy: We measure how well each method predicts whether a model answers an item correctly. These metrics include accuracy, F1 score, and ROC AUC.
  • Rank stability: We assess the consistency of model rankings by splitting the test set into two subsets. For each subset, we estimate model abilities and generate a ranking. Kendall’s $\tau$ is then computed between the two rankings.
4.2 Estimation Ability of PSN-IRT
Table 2 shows the estimation accuracy and reliability of PSN-IRT compared to traditional IRT and Deep-IRT. Both Deep-IRT and PSN-IRT significantly outperform traditional IRT in prediction accuracy. Despite its simpler Logistic output layer rooted in the IRT formula, PSN-IRT attains prediction accuracy on par with the more complex Deep-IRT, and surpasses it in rank reliability.
Given our primary objective of analyzing test sets, all methods except Deep-IRT utilized the more comprehensive 4PL structure for modeling benchmark items. However, the experimental results indicate that employing different parameter structures has a minimal impact on prediction accuracy, with detailed results available in Appendix A.3.
Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher Total
Chinese SimpleQA 2 3 1 9 3 5 21
MBPP 5 6 4 8 4 4 26
MATH 4 1 2 7 7 10 27
TheoremQA 1 10 3 11 2 2 27
GPQA Diamond 3 11 6 10 1 1 30
MMLU 9 9 9 5 5 3 31
BBH 6 5 8 6 6 7 32
GSM8K 8 2 5 3 11 11 32
HellaSwag 10 7 10 2 8 6 33
HumanEval 7 4 7 4 9 9 33
ARC-C 11 8 11 1 10 8 38
Table 3: Ranks of datasets based on average item-level properties estimated by PSN-IRT. Each dataset’s score is the mean of its benchmark items’ property. Lower values indicate better performance in that property.
Figure 3: Distribution of item-level properties across 11 LLM benchmarks.
5 Benchmark Analysis with PSN-IRT
5.1 Benchmark Quality Measurements
In this section, we delve into a detailed analysis of the benchmarks based on different measurements. Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) Vania et al. (2021) and Fisher information Lord (1980).
LEH score assesses the potential of a test example to evaluate future progress in LLMs. It is calculated as the derivative of the item’s Item Characteristic Curve (ICC) with respect to the highest observed latent ability. A high LEH score indicates that even the top model remains far from the saturation point of the ICC.
Fisher information measures the amount of information an item provides about a model’s ability level. Items with higher Fisher information are more informative for estimating model abilities. For 4PL IRT, it is defined as:
$I(\theta)=\frac{a^{2}(P(\theta)-c)^{2}(d-P(\theta))^{2}}{(d-c)^{2}P(\theta)(1-P(\theta))},$ (4)
where $P(\theta)$ is defined as Equation 2.
Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
5.2 Benchmark Analysis
Based on the measurements, we present an aggregate rank for each benchmark based on the sum of individual ranks in Table 3, where a lower total score indicates better overall quality. We also present the distribution of item-level properties across 11 LLM benchmarks in Figure 3. Moreover, we further illuminates the relationship between item difficulty and discriminability in Figure 4. Based on these experiment results, we present our detailed analysis as follows.
Finding 1: LLM benchmarks fail to achieve simultaneous excellence across multiple measurement properties.
As shown in Table 3, the results clearly demonstrate this lack of simultaneous excellence. While Chinese SimpleQA achieves the best overall score, no benchmark performs universally well across the individual metrics. For instance, Chinese SimpleQA itself suffers from poor feasibility with a rank of 9. This pattern of strengths being offset by significant weaknesses is common across benchmarks, which underscores the inherent challenges and trade-offs in benchmark design. Moreover, the property distributions visualized in Figure 3 further highlight this variability and the difficulty for any single benchmark to achieve balanced, high quality across all desired measurement properties.
Finding 2: Current LLM benchmarks exhibit an insufficient difficulty ceiling to challenge the most advanced models.
As shown in Table 3, difficulty is varied among different benchmarks. Datasets such as TheoremQA, Chinese SimpleQA, and GPQA Diamond lead in difficulty and continue to challenge many LLMs. In contrast, other benchmarks including ARC-C, HellaSwag, and HumanEval have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
Crucially, as shown in Figure 3, even within more challenging benchmarks, items often present only moderate difficulty to elite LLMs. For instance, item difficulty scores in the IRT scaling may be as low as 1.0 for models exhibiting ability scores above 3.0. This suggests that while such benchmarks can assess average models, they are less effective at separating most advanced systems, suggesting that they are insufficient for comprehensively evaluating frontier model capabilities.
Finding 3: Generally low LEH scores across most benchmarks reveal widespread item saturation.
While Table 3 indicates that datasets like GPQA Diamond and TheoremQA achieve the highest relative ranks for average LEH, suggesting they offer more headroom than other benchmarks, the absolute LEH values in Figure 3 are often smaller than ideal. Conversely, benchmarks such as MATH and particularly GSM8K exhibit even lower average LEH scores. This signifies that current high-performing models are already approaching or have reached the performance ceiling for a substantial portion of items within these test sets. This necessitates the development of new benchmark items and datasets explicitly designed with greater headroom as models continue to evolve.
Finding 4: Outlier-high guessing-rates in many benchmark items flag risks of data contamination.
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data Zhuang et al. (2024).
Finding 5: Widespread low-feasibility items in scientific benchmarks reveal flawed question design.
Feasibility estimates how well a question can be answered based on the information provided. Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear or defensible answer. As shown in Figure 3, datasets like TheoremQA and GPQA Diamond exhibit lower feasibility, likely due to complex scientific contexts or vague problem descriptions. In contrast, ARC-C and HellaSwag have high feasibility, suggesting their questions are well-formed and specific. Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities Rodriguez et al. (2021).
Finding 6: Item discriminability is weakened by excessively high or low item difficulty.
As shown in Table 3, benchmarks such as MATH and GSM8K achieve high discriminability, while GPQA Diamond and TheoremQA exhibit the lowest discriminability. This limited variability in model outcomes on these latter datasets may arise from too high item difficulty, thereby reducing their capacity for precise model differentiation.
Moreover, as shown in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—being either too easy or too hard. Consequently, datasets that feature a balanced distribution of difficulty often exhibit higher discriminability, enabling them to more effectively distinguish between models of varying capabilities.
6 Efficient Benchmarking with PSN-IRT
The pursuit of comprehensive LLM evaluation has often led to the creation of increasingly large and diverse benchmark datasets. However, the sheer volume of benchmark items does not inherently guarantee higher quality or more insightful evaluations, and can incur significant computational costs. In this section, we explore how PSN-IRT can facilitate more efficient benchmarking by guiding the strategic curation of smaller, yet highly informative, item sets that maintain or even enhance evaluation accuracy and reliability compared to larger, undifferentiated collections.
To empirically examine this principle, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involve Benchmark Agreement Testing (BAT) Perlitz et al. (2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena Chiang et al. (2024) and OpenCompass Arena Contributors (2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities.. Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess differentiation among stronger contenders.
As shown in Table 4, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking. For instance, a carefully selected subset composed of just 1000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048. This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
In contrast, other item selection metrics such as discriminability proves less effective for differentiating among strong models. While such items are typically adept at distinguishing between broad capability tiers (e.g., strong versus weak models), they often lack the granularity to capture nuanced performance differences at the higher end of the ability spectrum. This further verify the estimation reliability of our proposed PSN-IRT.
Metric w/ weak models w/o weak models
Variance Kendall Variance Kendall
All 0.0434 0.6444 0.0010 0.2381
Top 400
Random 0.0421 0.6444 0.0007 0.2381
Discrim. 0.1841 0.5556 0.0000 0.2381
Fisher 0.0049 0.6889 0.0033 0.7143
Top 1000
Random 0.0406 0.6444 0.0009 0.2381
Discrim. 0.1805 0.6000 0.0000 0.3813
Fisher 0.0039 0.6889 0.0030 0.9048
Table 4: Variance and Kendall’s $\tau$ for various data selection strategies with and without weak models.
7 Conclusion
In this work, we introduce PSN-IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks. Through extensive experiments on 12 LLMs across 11 diverse datasets, we demonstrate that current benchmarks often suffer from uneven measurement properties, insufficient difficulty ceilings, item saturation, and data contamination. Moreover, we show that strategically selecting smaller, high-quality item sets using PSN-IRT enhances alignment with human preference and separability among top models.
Limitations
Our work still has some limitations: 1) The current analysis is based on a specific selection of contemporary LLMs and benchmarks. Future studies could extend our analysis to encompass a broader and more diverse range of models and datasets as they emerge. 2) PSN-IRT is designed to process binary (correct/incorrect) outcome data. This restricts its direct applicability to benchmarks that utilize graded scoring Zheng et al. (2023), for which extensions incorporating more generalized psychometric models would be necessary. 3) The item parameters estimated by PSN-IRT are currently dependent on the items being included in the training data, meaning that analyzing entirely new, unseen items typically requires retraining the model. Future iterations could explore the integration of item content features to potentially enable parameter estimation for new items without full retraining.
Appendix A Appendix
A.1 Detailed Statistics
We provide the full list of average item-level property estimates for each dataset in Table 5. These values are computed as the mean of all item parameters within a benchmark and include difficulty, discriminability, guessing-rate, feasibility, LEH score, and Fisher information. The table allows for fine-grained comparisons between benchmarks beyond their aggregated rankings.
Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher
ARC-C -1.1637 1.2822 0.2621 0.9481 0.0041 0.0005
BBH -0.6033 1.3845 0.1657 0.8747 0.0057 0.0006
Chinese SimpleQA 0.2656 1.4474 0.0609 0.7088 0.0089 0.0006
GPQA Diamond 0.1902 0.8598 0.1518 0.6437 0.0142 0.0014
GSM8K -0.7899 1.5255 0.1491 0.9167 0.0038 0.0004
HellaSwag -1.0353 1.2920 0.2449 0.9304 0.0047 0.0006
HumanEval -0.6351 1.4459 0.1565 0.8957 0.0046 0.0005
MATH -0.2396 1.5793 0.0673 0.8290 0.0054 0.0004
MBPP -0.3456 1.3625 0.1345 0.8221 0.0071 0.0007
MMLU -0.8383 1.1968 0.2381 0.8921 0.0067 0.0008
TheoremQA 0.6548 1.0656 0.0888 0.5012 0.0145 0.0013
Table 5: Average item-level properties across datasets. Each value represents the mean of the corresponding metric over all benchmark items.
A.2 Case Study
Figure 5: Examples of benchmark items exhibiting high and low estimated values for key psychometric parameters, as determined by PSN-IRT. Each pair shows an item with a high parameter value alongside one with a low value for the specified characteristic.
As depicted in Figure 5, these parameters capture a wide range of item behaviors:
  • For difficulty, the figure contrasts a complex coding task with a simple factual recall question, showcasing items at opposite ends of the challenge spectrum.
  • Examples for discriminability include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.
  • For guessing-rate, an alphabetical sorting task exhibits a very low estimated parameter, offering little opportunity for chance success. In contrast, a multiple-choice question concerning scientific studies shows an exceptionally high guessing parameter. This item’s content is publicly accessible online, suggesting an associated risk of data contamination.
  • For feasibility, a logical arrangement task shows high feasibility, implying clarity and high attainability for proficient models, while a domain-specific question exhibits low feasibility, suggesting potential item ambiguities.
A.3 Ablation Study
To investigate the impact of model configurations, we evaluate PSN-IRT and baseline methods across 1PL to 4PL parameters, as shown in Table 6. Among these models, the Two-Parameter Logistic (2PL) model builds upon the 1PL by incorporating an item discriminability parameter $a$:
$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (5)
Subsequently, the Three-Parameter Logistic (3PL) model extends the 2PL by adding a guessing-rate parameter $c$:
$P(X=1\mid\theta,a,b,c)=c+(1-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (6)
The results demonstrate that performance exhibits no consistent pattern with increasing parameter complexity. While PSN-IRT with 4PL achieves the highest performance, simpler models occasionally outperform more complex ones in traditional IRT methods.
Additionally, we explore replacing one-hot encodings with learned embeddings and substituting the multi-layer perceptron with a graph neural network to model LLM-item interactions. However, these modifications failed to yield significant improvements, likely due to the inherent simplicity of LLM-item interaction patterns that don’t require complex relational modeling. We will explore improved input representations and model architectures as future work to enhance model robustness.
Model Method Parameter ACC F1 AUC Kendall Average
IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000 0.8056
2PL 0.7172 0.7998 0.6963 0.9091 0.7806
3PL 0.7218 0.8031 0.7057 0.9697 0.8001
4PL 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 1PL 0.7168 0.8011 0.7268 0.9697 0.8036
2PL 0.7232 0.8130 0.7301 0.9697 0.8090
3PL 0.7172 0.8081 0.7297 1.0000 0.8138
4PL 0.7070 0.7811 0.7278 0.9697 0.7964
VI 1PL 0.7209 0.8024 0.6822 0.9697 0.7938
2PL 0.7198 0.8012 0.6938 1.0000 0.8037
3PL 0.7189 0.8011 0.6956 0.9697 0.7963
4PL 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 1PL 0.7007 0.7801 0.6901 0.9697 0.7852
2PL 0.7301 0.8048 0.6987 0.8788 0.7781
3PL 0.7172 0.7993 0.7048 0.9091 0.7826
4PL 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT Deep Learning 1PL 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091 0.8502
2PL 0.7560 0.8266 0.8259 0.9697 0.8445
3PL 0.7965 0.8510 0.8510 0.8788 0.8443
4PL 0.7991 0.8520 0.8477 1.0000 0.8747
Table 6: Comparison of prediction accuracy and rank reliability across different methods.